OpenStudy (anonymous):

Trig.. trying to determine how many triangles and find the missing sides/ angles... b=5, c=3, B = 100 .. Help please

Directrix (directrix):

This is the Ambiguous Case in solving triangles because you are given two sides and the non-included angle and asked to find the other angles and sides of the triangle. Using the Law of Sines, Sine C /3 = Sine 100/5 Sine C / 3 = .985 / 5 Sine C = [.985 (3)] / 5 Sine C = .5908 C = 36.22 degrees OR 180 - 36.22 = 143.78 degrees If C were 143.78 degrees, it and the 100 degrees of angle B could not be in the same triangle. Therefore, C = 36.22 degrees is the only viable option for C and there is only one possible triangle. In that one triangle, angle A has measure 43.78 degrees [180 - 36.22 - 100] Continuing with the Law of Sines, Sine A / a = = Sine 100 / 5 Sine 43.78 / a = .985 / 5 .692 / a = .985 / 5 .985a = 5 ( .692 ) a = 3.513 Angles Sides A: 43.78° a = 3.513 B: 100° b = 5 C: 36.22° c = 3